OpenStudy (anonymous):

A box has a weight of 120 lbs and the bottom of the box is 12 in2 . What is the pressure the box exerts on the floor?

OpenStudy (anonymous):

Pressure is defined as P=F/A where F= force and A=area of effect in this case the only force you can recall would be weight. Weight is W=m*g where g=gravitational acceleration, 32.17 ft/s^2 So... P=m*g/A You know m=120lbs A=12 in^2 and g can be considered a constant So you can solve for Pressure

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

ok im lost again , so i divide 12in into120lbs ?

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

the area, A=12 in^2 the mass, m=120 lbs the gravitational constant, g=32 ft/s^2=384 in/s^2 ft=feet in=inches s=seconds lbs=pounds P=pressure \[P=\dfrac {m \times g} {A}\] \[P=\dfrac {120lbs \times 384 ( inches/s^2)} {12 inches^2} \]

OpenStudy (anonymous):

Thank You again

OpenStudy (anonymous):

by the way, pounds are a mass unit, not a weight unit.

OpenStudy (anonymous):

A bronze statue weighs 2400newtons & has a base that is 4 meters 1/2 meters. what is the pressure statue exerts on the floor ?

OpenStudy (anonymous):

 That's your base, so the area of effect would be the area of that quadrilateral Now, they give you the Force (Weight) that is 2400 N N=Newtons Now, N/m^2 = Pa Pa= Pascal, a presure unit We recall the pressure eq P=F/A And you know the force, 2400N You can calculate the area (area of a quadrilateral) And then you'll obtain the pressure

OpenStudy (anonymous):

ughhhh , so how my question going to look when i solve it ?